OpenStudy (anonymous):

What is the convergent sum of the series? Series is: 640, 320, 160, 80, 40.........

OpenStudy (anonymous):

The sequence is geometric, since each term is the previous term multiplied by \(\frac{1}{2}\). The series is \[ \sum_{n=0}^{\infty} 640\left(\frac{1}{2} \right)^n.\] You can find the sum to infinity of a geometric series using the formula \[ \sum_{n=0}^{\infty} ar^n = \frac{a}{1-r}\] as long as -1<r<1. (In this case, a=640, and r=1/2.)

OpenStudy (anonymous):

it is a G.P a=640,r=320/640=1/2 \[S \infty =\frac{ a }{ 1-r }=\frac{ 640 }{1-\frac{ 1 }{ 2 } }=\frac{ 640 }{\frac{ 1 }{ 2 } }\] =640*2=1280

OpenStudy (anonymous):

Thanks guys.

OpenStudy (anonymous):

yw